OpenStudy (anonymous):

a carbon oxygen double bond in a certain organic molecule absorbs radiation that has a frequency of 6.0 *10^13 s^-1 what is the wavelength of this radiation Energy of the ratdiation per photon per mole of photons

OpenStudy (anonymous):

wavelength = speed of light/frequency you're given frequency, so just plug into equation to find wavelength Energy = plank's constant x frequency so you are given frequency here, plug into equation to find energy

OpenStudy (anonymous):

so \[\lambda=c/v?\] and E=H * V

OpenStudy (anonymous):

correct

OpenStudy (anonymous):

What does s^-1 stand for?

OpenStudy (anonymous):

that's the unit

OpenStudy (anonymous):

which is hertz per second?

OpenStudy (anonymous):

just hertz, not hertz per second

OpenStudy (anonymous):

thank you